Please help me write the equation for the following relations.
R = {(x, y): (4, 5), (8, 7), (12, 9), (16, 11), . . .}
R = {(x, y): (2, 3), (4, 4), (6, 5), (8, 6), . . .}
Q = {(x, y): (2, 8), (3, 27), (4, 64), (5, 125), . . .}
P = {(x, y): (1, 0), (2, 4), (3, 8), (4, 12), . . .}

Answer by stanbon(75887)

Please help me write the equation for the following relations.
R = {(x, y): (4, 5), (8, 7), (12, 9), (16, 11), . . .}
Notice that y increases by 2 when x increases by 4
slope = 2/4 = 1/2
Notice that if you decrease x by 4 y will decrease by 2
So, y-intercept = (0,3)
Equation: y = (1/2)x+3
---------------------
R = {(x, y): (2, 3), (4, 4), (6, 5), (8, 6), . . .}
slope = 1/2
y-int: (0,1)
Equation: y = (1/2)x + 1
---------------------
Q = {(x, y): (2, 8), (3, 27), (4, 64), (5, 125), . . .}
Not a linear relation because y does not increase consistently
as x increases.
Equation: y = x^3
------------------------------
P = {(x, y): (1, 0), (2, 4), (3, 8), (4, 12), . . .}
slope = 4
y-int: -4
Equation: y = 4x-4
----------------------
Cheers,
Stan H.
================